OpenStudy (anonymous):

if y varies directly as x and inversely as x and y=-18 and z=3 when x=6

OpenStudy (anonymous):

ok so i'll give you what my dad gave me y varies directly with x y=kx 10=k(15) k=10/15 y=(10/15)x when x=6 y=(10/15)(6)=60/15=4 x varies inversely as y x=k/y 2=k/8 k=16 x=16/y when y=17 x=16/17 now you figure the rest out oh and I didn't use the numbers you gave you need to do that yourself or you wont learn anything